# Deep Learning for Continuous-Time Stochastic Control with Jumps

### v2

@@ Appendix C Linear-quadratic regulator with jumps: detailed results
The value function for the LQR problem with jumps (25) is given by
@@
$0=\partial_{t}V(t,x)+\frac{1}{2}\text{Tr}\left[\Sigma\Sigma^{\top}\hskip-0.85358pt\nabla^{2}V_{x}(t,x)\right]\hskip-1.13809pt+\inf_{a\in\mathbb{R}^{d}}\Big\{c_{1}\|a\|^{2}_{2}$ (29)
$\hskip 85.35826pt+(\Lambda_{1}+\Lambda_{2}\|a\|^{2}_{2})\,\mathbb{E}\left[V(t,x+Z_{1})-V(t,x)\right]+a^{\top}\hskip-0.56905pt\nabla_{x}V(t,x)\Big\}$
with terminal condition $V(T,x)=c_{2}\|x\|^{2}_{2}$. It is straightforward to show using the verification Theorem 2.2 with HJB equation (29) and the Ansatz
$V(t,x)=\frac{1}{2}h(t)\|x\|^{2}_{2}+g(t)\,$, (30)
that $g$ satisfies
$g(t)=\frac{1}{2}\big(\text{Tr}[\Sigma\Sigma^{\top}]+\Lambda_{1}\hskip 0.28453pt\upsilon)\int_{t}^{T}h(s)ds\,$, (31)
where $h$ solves the non-linear ordinary differential equation
$h^{\prime}(t)=\frac{h^{2}(t)}{2c_{1}+h(t)\hskip 0.56905pt\upsilon\hskip 0.56905pt\Lambda_{2}},\quad h(T)=2c_{2}\,$, (32)
@@
which can be computed efficiently using a numerical method such Runge–Kutta of order $5(4)$ (Dormand & Prince, 1980). The optimal control $\alpha^{*}$ is then given by
$\alpha^{*}(t,x)=-\frac{h(t)x}{2c_{1}+h(t)\hskip 0.56905pt\upsilon\hskip 0.28453pt\Lambda_{2}}\,.$ (33)
@@
In the constant intensity case $\Lambda_{2}=0,\Lambda_{1}>0$, (31) and (32) simplify to
$h(t)=\frac{2c_{1}c_{2}}{c_{1}+c_{2}(T-t)}\quad\text{and}\quad g(t)=(\text{Tr}[\Sigma\Sigma^{T}]+\Lambda_{1}v)\,c_{1}\left(\ln(\frac{c_{2}}{c_{1}}(T-t)+1)\right)\,.$
In this section, we assume for each jump size $Z_{j}\sim\mathcal{N}_{d}(0_{d},\Sigma_{J})$, with $\Sigma_{J}\in\mathbb{R}^{d\times d}$. Hence, we sample according to $\mathcal{Z}$ being a Gaussian distribution the $M_{3}$ points $z_{j}^{(m)}$ of GPI-PINN for each $(t_{m},x_{m})$, $m=1,\ldots,M_{1}$, and the $M_{1}$ points $z_{m}$ of GPI-CBU, see Section 4. Moreover, the frequencies with which various states appear in the sample should be roughly proportional to the probabilities of their occurrence under an optimal policy (Bachouch et al., 2022). As one can guess from (28), it is optimal to drive the process $X^{\alpha}$ towards 0. Therefore, since $\mu(t,x)=\mu(x\mid t)\mu(t)$, we assume $\mu(t)\sim\mathcal{U}_{[0,T]}$ and $\mu(x\mid t)\sim\sqrt{t}\,\mathcal{N}_{d}(0_{d},I_{d})$ to generate the values of $x$ and $t$. Similarly, $\nu(x)\sim\sqrt{T}\,\mathcal{N}_{d}(0_{d},I_{d})$. We then use as parameters of the LQR: $T=1$, $c_{1}=1$, $c_{2}=0.25$ and each element of $\Sigma$, $\Sigma_{J}$ is sampled from $\mathcal{U}_{[0,1]}$.
The three following Figures 7(a)-7-8 are obtained with GPI-CBU in $d=50$ dimensions under the constant intensity case $\Lambda_{1}=0.25$, $\Lambda_{2}=0$.
@@
(a) Value function $V(t,x)$ (left) and first component of the optimal control $\alpha^{*}_{1}(t,x)$ (right) at $t=0$ for $x=(x_{1},0,\dots,0)$ with $x_{1}\in[-2.5,2.5]$. Orange dotted line: numerical results of GPI-CBU with $\pm 1$ standard deviation given by orange shaded area ($k_{*}=1500$, $\Lambda_{2}=0$). Blue line: analytical solution (30)-(33).
@@
Figure 7: Value function $V(t,x)$ (left) and first component of the optimal control $\alpha^{*}_{1}(t,x)$ (right) at $x=\mathbf{1}_{50}$ with $t\in[0,1]$. Orange dotted line: numerical results from GPI-CBU with $\pm$ one standard deviation in orange shaded area ($k_{*}=1500$, $\Lambda_{2}=0$). Blue line: analytical solution (30)-(33).
@@
Figure 8: Heatmaps of ${\rm MAE}_{V}$ (left) and ${\rm MAE}_{\alpha}$ (right) for $t\in[0,1]$ and for $x=(x_{1},\bm{1}_{49})$ with $x_{1}\in[-2.5,2.5]$, obtained from GPI-CBU for the LQR problem with $\Lambda_{2}=0$.
We now consider in the following results the controlled intensity case with $\Lambda_{1}=0,\Lambda_{2}=2$, using the same parameters and sampling procedure as above. Since the intensity now also depends on the control, GPI-PINN becomes even more computationally intensive than in the constant intensity case and of no practical interest in the presence of jumps. Consequently, in the following Figures 910 (and 2), we again focus exclusively on GPI-CBU for a $50$-dimensional LQR problem.
@@
Figure 9: Value function $V(t,x)$ (left) and first component of the optimal control $\alpha^{*}_{1}(t,x)$ (right) at $x=\mathbf{1}_{50}$ with $t\in[0,1]$. Orange dotted line: numerical results from GPI-CBU with $\pm$ one standard deviation in orange shaded area ($k_{*}=1500$, $\Lambda_{1}=0$). Blue line: analytical solution (30)–(33).
@@
Figure 10: Heatmaps of ${\rm MAE}_{V}$ (left) and ${\rm MAE}_{\alpha}$ (right) for $t\in[0,1]$ and for $x=(x_{1},\bm{1}_{49})$ with $x_{1}\in[-2.5,2.5]$, obtained from GPI-CBU for the LQR problem with $\Lambda_{1}=0$.
These results further validate the accuracy of GPI-CBU, even when considering a controlled intensity for the jump process. As expected, the optimal control $\alpha_{1}^{*}$ is lower in magnitude compared to the constant intensity setting, since it increases the intensity and thus the likelihood of jumps, while the optimal strategy still consists in driving $X^{\alpha}$ towards 0. Again, higher ${\rm MAE}_{V}$ and ${\rm MAE}_{\alpha}$ metrics, computed from (30)–(33), tend to occur at the boundaries of the domain, see Figure 10.
Finally, you can find below a table summarizing higher-dimensional results for GPI-CBU in the LQR setting with controlled intensity after $k_{*}=5000$ epochs. This illustrates the scalability of the approach.
Dimensions $d$ $MAE_{V}$ $MAE_{\alpha}$ Loss $\mathscr{L}_{1}$ Loss $\mathscr{L}_{2}$ Time (sec.)
@@
Table 1: GPI-CBU’s performance metrics in function of the state dimension $d$ for the LQR example with controlled intensity after $k_{*}=5000$ epochs.
@@ D.1 Constant coefficients
We first study the optimal consumption-investment problem with constant coefficients, where we assume the risk-free asset evolves according to
@@
for a constant interest rate $r\in\mathbb{R}$, and with $n$ stocks whose price follows for $i=1,\ldots,n$,
$\frac{dS^{i}_{t}}{S^{i}_{t-}}=\mu^{i}dt+\Sigma^{\top}_{i}dW_{t}+d\sum_{j=1}^{M^{i}_{t}}\left(e^{Z^{i}_{j}}-1\right)$ (35)
for some constants $\mu^{i}\in\mathbb{R}$, $\Sigma_{i}\in\mathbb{R}^{n}$, correlated [^8] Note that $\Sigma=(\Sigma_{1},\ldots,\Sigma_{n})^{\top}\hskip-2.27621pt\in\mathbb{R}^{n\times n}$ can be seen as the upper Choleski decomposition of the correlation matrix $\Sigma\Sigma^{\top}$ of the Brownian motion $W$. $n$-dimensional Brownian motions $W$, independent Poisson processes $M^{i}$ with constant intensities $\lambda^{i}\geq 0$ and i.i.d. normal random vectors $Z^{i}_{1},Z^{i}_{2},...,Z^{i}_{M^{i}_{t}}\sim\mathcal{N}(\mu^{i}_{Z},\sigma^{i}_{Z})$ with $\mu^{i}_{Z}\in\mathbb{R},\sigma^{i}_{Z}\in\mathbb{R}^{+}$. Suppose an investor starts with an initial endowment of $Y_{0}>0$, consumes at rate $c_{t}Y^{\alpha}_{t-}$ and invests in the $n$ stocks according to $\pi^{i}_{t}Y^{\alpha}_{t-}$, $i=1,\dots n$. If the risk-free asset is used to balance the transactions, the resulting wealth evolves according to
$\frac{dY^{\alpha}_{t}}{Y^{\alpha}_{t-}}=\left(r+\sum_{i=1}^{n}\pi^{i}_{t}(\mu^{i}-r)-c_{t}\right)dt+\sum_{i=1}^{n}\pi^{i}_{t}\!\left[\Sigma^{\top}_{i}dW_{t}+d\sum_{j=1}^{M^{i}_{t}}\left(e^{Z^{i}_{j}}-1\right)\right].$ (36)
Let us assume she attempts to optimize
$\mathbb{E}\!\left[\int_{0}^{T}e^{-\rho s}u(c_{s}Y^{\alpha}_{s})ds+U(Y^{\alpha}_{T})\right]\,$,
for two utility functions $u,U\colon\mathbb{R}^{+}\to\mathbb{R}$ and $\rho\in\mathbb{R}$. Since the driving noise in (36) has stationary independent increments, it is enough to consider strategies of the form $c_{t}=c(t,Y^{\alpha}_{t-})$ and $\pi^{i}_{t}=\pi^{i}(t,Y^{\alpha}_{t-})$ for functions $c,\pi^{i}\colon[0,T)\times\mathbb{R}^{+}\to(0,1)$, $i=1,\dots,n$. This $n+1$-dimensional control is then written $\alpha_{t}=(\pi^{1}_{t},\ldots,\pi^{n}_{t},c_{t})^{\top}\hskip-2.84526pt\in A:=(0,1)^{n+1}$. Assuming constant relative risk aversion (CRRA) utility functions $u,U$ with $\gamma\in(0,1)$, the reward functional is then denoted
$V^{\alpha}(t,y)=\mathbb{E}\left[\int_{t}^{T}e^{-\rho(s-t)}\frac{\left(c_{s}Y^{\alpha}_{s}\right)^{\gamma}}{\gamma}ds+\frac{\left(Y^{\alpha}_{T}\right)^{\gamma}}{\gamma}\,\Big|\,Y^{\alpha}_{t}=y\right].$ (37)
Finally, writing $\mu=(\mu^{1},\ldots,\mu^{n})^{\top}$, $\mu_{Z}=(\mu_{Z}^{1},\ldots,\mu_{Z}^{n})^{\top}$, $\sigma_{Z}=(\sigma^{1}_{Z},\ldots,\sigma^{n}_{Z})^{\top}$, $\lambda=(\lambda^{1},\ldots,\lambda^{n})^{\top}$, $\pi=(\pi^{1},\ldots,\pi^{n})^{\top}\hskip-1.42262pt$, the associated HJB equation for the value function $V(t,y)=\sup_{\alpha}V^{\alpha}(t,y)$ satisfies [^9] Note that the term $\rho V(t,y)$ in (38) is not contained in our general HJB equation (4). But GPI-PINN and GPI-CBU still work if it is added. for all $(t,y)\in[0,T)\times\mathbb{R}^{+}$,
@@
$\hskip-6.25958pt+\frac{1}{2}\pi^{\top}\hskip-1.42262pt\Sigma\Sigma^{\top}\hskip-1.42262pt\pi\,y^{2}\,\partial^{2}_{yy}V(t,y)+\sum_{i=1}^{n}\lambda^{i}\mathbb{E}\left[V(t,y+y\,\pi^{i}(e^{Z^{i}_{1}}-1))-V(t,y)\right]+\frac{(cy)^{\gamma}}{\gamma}\bigg\}\,$,
@@
with $V(T,y)=y^{\gamma}/\gamma$. This stochastic control problem does not admit an analytical solution but we can instead characterize its solution in terms of a PIDE, so as to assess the accuracy of the proposed Algorithms 1 and 2. Following Øksendal & Sulem (2007), we assume the following form $V(t,y)=A(t)y^{\gamma}/\gamma$ for the value function. First order condition on the HJB equation (38) then implies that optimal fractions of wealth is a vector of constants $\pi^{*}(t,y)=\pi^{*}:=(\pi^{1,*},\ldots,\pi^{n,*})^{\top}$ satisfying
$(\mu-r)^{\top}+\Sigma\Sigma^{\top}\pi^{*}(\gamma-1)+\sum_{i=1}^{n}\lambda^{i}\,\mathbb{E}\left[\left(1+\pi^{i,*}(e^{Z^{i}_{1}}-1)\right)^{\gamma-1}\left(e^{Z^{i}_{1}}-1\right)\right]=0\,$, (39)
and the optimal consumption rate $c^{*}$ is independent from the wealth $y$ and given by
$c^{*}(t,y)=A(t)^{1/(\gamma-1)}\,.$ (40)
@@
Plugging back these optimal controls inside the HJB equation (38), we find that the function $A(t)$ satisfies the ODE
$\frac{\partial_{t}A(t)}{A(t)}=\rho\hskip 1.42262pt$- $\gamma\left(r+(\mu-r)^{\top}\pi^{*}-A(t)^{1/(\gamma-1)}\right)-\frac{1}{2}\gamma(\gamma-1)\pi^{*\top}\Sigma\Sigma^{\top}\pi^{*}$ (41)
$\hskip 59.75078pt-\sum_{i=1}^{n}\lambda^{i}\mathbb{E}\left[\left(1+\pi^{i,*}(e^{Z^{i}_{1}}-1)\right)^{\gamma}\hskip-1.42262pt-1\right]-A(t)^{1/(\gamma-1)},$
@@
with $A(T)=1$, which can be solved numerically using again the Runge–Kutta method of order $5(4)$. For both Algorithms 1 and 2, we sample the time and space points independently from the uniform distributions $\mathcal{U}_{[0,T]}$ and $\mathcal{U}_{[0,y_{b}]}$, respectively. The parameters of the optimal investment problem below are as follows: $T=1,y_{b}=150,r=0.02,\rho=0.045,\gamma=0.3,\lambda=0.45\cdot\mathbf{1}_{n},\mu_{Z}=0.25\cdot\mathbf{1}_{n},\sigma_{Z}=0.2\cdot\mathbf{1}_{n},\mu=0.032\cdot\mathbf{1}_{n},\Sigma\Sigma^{\top}=0.2\cdot\mathbf{1}_{n\times n}$ with $\text{diag}(\Sigma\Sigma^{\top})=\mathbf{1}_{n}$. Note that, compared to the LQR problem in Section 5.1, the proportionality factor $\xi$ is now set to $10$. Moreover, since $A=(0,1)^{n+1}$, $\sigma_{\alpha_{out}}$ is chosen to be the sigmoid activation function. Instead of the DGM architecture, we here train a classical feedforward neural network with 4 hidden layers, each of 128 neurons. Finally, the ${\rm MAE}_{V}$ and ${\rm MAE}_{\alpha}$ metrics are again defined as in Section 5.1 on a test set of size $M$, uniformly sampled from $0,1]\times[0,y_{b}]$ with $V$ and $\alpha^{*}$ obtained from the RK45 method described above.
@@
We again compare in Figure 11 the ${\rm MAE}_{V}$ and runtime between GPI-PINN and GPI-CBU in function of the number of epochs $k$ for $n=10$ stocks in the consumption-investment problem (37) with and without jumps. We again see that the residual-based GPI-PINN Algorithm 1 tends to be more stable and accurate for larger epochs, despite having a higher runtime. When accounting for jumps in the dynamics (36), GPI-PINN becomes numerically very time-consuming, even for a 10-dimensional control problem. This issue is even amplified compared with the LQR problem (see Figure 2(a)), as the jump size now depends on the control $\pi$ in the wealth dynamics (36). In contrast, GPI-CBU again handles these jumps far more efficiently.
@@
Figure 11: Comparison of the ${\rm MAE}_{V}$ (blue) and the runtime in seconds (green) between GPI-PINN (solid line) and GPI-CBU (dashed line) in function of the number of epochs $k$ for the optimal consumption-investment problem without jumps (left) and with jumps (right) for $n=10$ stocks.

### v3

@@ Appendix C Linear-quadratic regulator with jumps: detailed results
The value function of the LQR problem with jumps (18) is given by
@@
$0=\partial_{t}V(t,x)+\frac{1}{2}\text{Tr}\left[BB^{\top}\hskip-0.85358pt\nabla^{2}V_{x}(t,x)\right]\hskip-1.13809pt+\inf_{a\in\mathbb{R}^{d}}\Big\{c_{1}\|a\|^{2}_{2}+a^{\top}\nabla_{x}V(t,x)$
$\hskip 85.35826pt+(\Lambda_{1}+\Lambda_{2}\|a\|^{2}_{2})\,\mathbb{E}\left[V(t,x+Z_{1})-V(t,x)\right]\Big\}$
with terminal condition $V(T,x)=c_{2}\|x\|^{2}_{2}$. Using the Ansatz
$V(t,x)=g(t)+\frac{1}{2}h(t)\|x\|^{2}_{2}\,$,
it is straightforward to see that $g$ is of the form
$g(t)=\frac{1}{2}\big(\text{Tr}[BB^{\top}]+\Lambda_{1}\hskip 0.28453pt\upsilon)\int_{t}^{T}h(s)ds$ (24)
for the unique solution of the non-linear ODE
$h^{\prime}(t)=\frac{h^{2}(t)}{2c_{1}+h(t)\hskip 0.56905pt\upsilon\hskip 0.56905pt\Lambda_{2}},\quad h(T)=2c_{2}\,.$ (25)
@@
(25) can efficiently be solved using a numerical method such as Runge–Kutta of order $5(4)$; see e.g. Dormand and Prince (1980). The optimal control $\alpha^{*}$ is given in terms of $h$ by
$\alpha^{*}(t,x)=-\frac{h(t)x}{2c_{1}+h(t)\hskip 0.56905pt\upsilon\hskip 0.28453pt\Lambda_{2}}\,.$
@@
Note that in the constant intensity case $\Lambda_{1}>0,\Lambda_{2}=0$, (24)–(25) admit the explicit solutions
$g(t)=(\text{Tr}[BB^{T}]+\Lambda_{1}v)\,c_{1}\left(\log(\frac{c_{2}}{c_{1}}(T-t)+1)\right)\quad\mbox{and}\quad h(t)=\frac{2c_{1}c_{2}}{c_{1}+c_{2}(T-t)}\,.$ (26)
But also in the case $\Lambda_{2}>0$, numerical solutions of (24)–(25) obtained with Runge–Kutta of order 5(4) provide very precise reference solutions $V$ and $\alpha^{*}$ to the LQR problem (18).
GPI-PINN and GPI-CBU both need random points $(t_{m},x_{m})\in[0,T)\times\mathbb{R}^{d}$ and $(y_{m})\in\mathbb{R}^{d}$ sampled from two distributions $\mu$ on $0,T)\times\mathbb{R}^{d}$ and $\nu$ on $\mathbb{R}^{d}$, which we chose according to Bachouch et al. (2022).
In the following we consider problem (18) with $T=1$, $B=I_{d}$, $c_{1}=1$, $c_{2}=1/4$ and $d$-dimensional standard normal jumps $Z_{j}$. Figures 7(a)8 show results obtained with GPI-CBU in $d=50$ dimensions for the constant jump intensity case $\Lambda_{1}=1/4$ and $\Lambda_{2}=0$, which admits analytical reference solutions.
@@
(a) Value function $V(0,x)$ (left) and first component of the optimal control $\alpha^{*}_{1}(0,x)$ (right) for $x=(x_{1},0,\dots,0)$ with $x_{1}\in[-2.5,2.5]$. Orange dotted lines: numerical results of GPI-CBU with $\pm 1$ standard deviation given by orange shaded area. Blue lines: analytical solution (26). $d=10$, $\Lambda=1/4$, $\Lambda_{2}=0$, $k_{*}=1500$.
@@
Figure 7: Value function $V(t,x)$ (left) and first component of the optimal control $\alpha^{*}_{1}(t,x)$ (right) for $t\in[0,1]$ and $x=\mathbf{1}_{50}$. Orange dotted lines: numerical results from GPI-CBU with $\pm 1$ standard deviation in orange shaded area. Blue lines: analytical solution (26). $d=50$, $\Lambda_{1}=1/4$, $\Lambda_{2}=0$, $k_{*}=1500$.
@@
Figure 8: Heatmaps of ${\rm MAE}_{V}$ (left) and ${\rm MAE}_{\alpha}$ (right) for $t\in[0,1]$ and $x=(x_{1},1,\dots,1)$ with $x_{1}\in[-2.5,2.5]$ obtained from GPI-CBU for the LQR problem with $d=50$, $\Lambda_{1}=1/4$ and $\Lambda_{2}=0$. $k_{*}=1500$.
${\rm MAE}_{\alpha}$ in Figure 8 is defined, analogously to ${\rm MAE}_{V}$ in Section 5.1, by
${\rm MAE}_{\alpha}=\frac{1}{M}\sum_{i=1}^{M}\|\alpha_{\phi^{(k_{*})}}\hskip-1.13809pt(t_{i},x_{i})-\alpha^{*}(t_{i},x_{i})\|_{2}\vskip-2.84526pt$
for a test set of size $M$ uniformly sampled from $0,1]\times[-2.5,2.5]^{d}$.
Next, we consider the same LQR problem (18) as above in 50 dimensions with a controlled jump intensity. Figures 910 (and 2), show results of GPI-CBU for $\Lambda_{1}=0$ and $\Lambda_{2}=2$. Reference solutions were computed with Runge–Kutta of order 5(4).
@@
Figure 9: Value function $V(t,x)$ (left) and first component of the optimal control $\alpha^{*}_{1}(t,x)$ (right) for $t\in[0,1]$ and $x=\mathbf{1}_{50}$. Orange dotted lines: numerical results of GPI-CBU with $\pm 1$ standard deviation in orange shaded area. Blue lines: reference solution. $d=50$, $\Lambda_{1}=0$, $\Lambda_{2}=2$, $k_{*}=1500$.
@@
Figure 10: Heatmaps of ${\rm MAE}_{V}$ (left) and ${\rm MAE}_{\alpha}$ (right) for $t\in[0,1]$ and $x=(x_{1},\bm{1}_{49})$ with $x_{1}\in[-2.5,2.5]$ obtained from GPI-CBU for the LQR problem with $d=50$, $\Lambda_{1}=1/4$ and $\Lambda_{2}=0$. $k_{*}=1500$.
These results illustrate the accuracy of GPI-CBU for high-dimensional control problems with controlled jumps. It can be seen that the optimal control $\alpha^{*}$ is lower in magnitude compared to the case of constant jump intensity since now, exerting control increases the likelihood of jumps. But in both cases, it is optimal to steer the state process $X^{\alpha}_{t}$ towards 0 as $t$ approaches the time horizon $T$.
Table 1 summarizes results of GPI-CBU for higher-dimensional LQR problems with the same parameters as above with controlled jump intensity $\Lambda_{1}=0$, $\Lambda_{2}=2$ obtained over $k_{*}=5000$ training epochs. This demonstrates the scalability of GPI-CBU.
Table 1: GPI-CBU performance metrics as a function of the state dimension $d$ for LQR problems of the form (18) with $\Lambda_{1}=0$, $\Lambda_{2}=2$ and $k_{*}=5000$ training epochs.
Dimensions $d$ ${\rm MAE}_{V}$ ${\rm MAE}_{\alpha}$ Loss $\widehat{\mathscr{L}}_{1}(\theta^{k_{*}})$ Loss $\widehat{\mathscr{L}}_{2}(\theta^{k_{*}})$ Time (sec)
@@ D.1 Constant coefficients
We first study the optimal consumption-investment problem in a model with constant coefficients, where the risk-free asset evolves according to
@@
for a constant interest rate $r\in\mathbb{R}$, and there are $n$ stocks with prices following
$\frac{dS^{i}_{t}}{S^{i}_{t-}}=\mu^{i}dt+\sigma^{i}dW^{i}_{t}+d\sum_{j=1}^{M^{i}_{t}}\left(e^{Z^{i}_{j}}-1\right),$ (28)
for constant drifts $\mu^{i}\in\mathbb{R}$ and volatilities $\sigma^{i}\in\mathbb{R}_{+}$, an $n$-dimensional Brownian motion $W$ with correlation matrix $\Sigma$, independent Poisson processes $M^{i}$ with constant intensities $\lambda^{i}\geq 0$ and i.i.d. normal random variables $Z^{i}_{1},Z^{i}_{2},...\sim\mathcal{N}(\mu^{i}_{Z},\sigma^{i}_{Z})$ with $\mu^{i}_{Z}\in\mathbb{R}$ and $\sigma^{i}_{Z}\in\mathbb{R}_{+}$. Suppose an investor starts with an initial endowment of $Y_{0}>0$, consumes at rate $c_{t}Y^{\alpha}_{t-}$ and invests $\pi^{i}_{t}Y^{\alpha}_{t-}$, $i=1,\dots n$, in the $n$ stocks. If the risk-free asset is used to balance the transactions, the resulting wealth evolves according to
$\frac{dY^{\alpha}_{t}}{Y^{\alpha}_{t-}}=\left(r+\sum_{i=1}^{n}\pi^{i}_{t}(\mu^{i}-r)-c_{t}\right)dt+\sum_{i=1}^{n}\pi^{i}_{t}\!\left[\sigma^{i}dW^{i}_{t}+d\sum_{j=1}^{M^{i}_{t}}\left(e^{Z^{i}_{j}}-1\right)\right].$ (29)
Let us assume the investor attempts to maximize
for two utility functions $u,U\colon\mathbb{R}_{+}\to\mathbb{R}$ and a discount factor $\rho\in\mathbb{R}$. Since the driving noise in (29) has stationary and independent increments, it is enough to consider strategies of the form $c_{t}=c(t,Y^{\alpha}_{t-})$ and $\pi^{i}_{t}=\pi^{i}(t,Y^{\alpha}_{t-})$ for functions $c,\pi^{i}\colon[0,T)\times\mathbb{R}_{+}\to\mathbb{R}_{+}$, $i=1,\dots,n$. We write this $n+1$-dimensional control as $\alpha_{t}=(c_{t},\pi^{1}_{t},\ldots,\pi^{n}_{t})^{\top}\hskip-2.84526pt\in A:=\mathbb{R}_{+}^{n+1}$. Assuming constant relative risk aversion (CRRA) utility functions $u,U$ with relative risk aversion $\gamma>0$ and denoting $\delta=1-\gamma$, the resulting reward functional is
$V^{\alpha}(t,y)=\mathbb{E}\left[\int_{t}^{T}e^{-\rho(s-t)}\frac{\left(c_{s}Y^{\alpha}_{s}\right)^{\delta}}{\delta}ds+e^{-\rho(T-t)}\frac{\left(Y^{\alpha}_{T}\right)^{\delta}}{\delta}\,\Big|\,Y^{\alpha}_{t}=y\right].$ (30)
Finally, writing $\mu=(\mu^{1},\ldots,\mu^{n})^{\top}$, $\sigma=\text{diag}(\sigma^{1},\ldots,\sigma^{n})$, $\mu_{Z}=(\mu_{Z}^{1},\ldots,\mu_{Z}^{n})^{\top}$, $\sigma_{Z}=(\sigma^{1}_{Z},\ldots,\sigma^{n}_{Z})^{\top}$, $\lambda=(\lambda^{1},\ldots,\lambda^{n})^{\top}$, $\pi=(\pi^{1},\ldots,\pi^{n})^{\top}\hskip-1.42262pt$, the associated HJB equation for the value function $V(t,y)=\sup_{\alpha}V^{\alpha}(t,y)$ satisfies [^8] Note that the term $\rho V(t,y)$ in (31) is not contained in our general HJB equation (3). But GPI-PINN and GPI-CBU still work if it is added. for all $(t,y)\in[0,T)\times\mathbb{R}_{+}$,
@@
$\hskip-6.25958pt+\frac{1}{2}\pi^{\top}\hskip-1.42262pt\sigma\Sigma\sigma^{\top}\hskip-1.42262pt\pi\,y^{2}\,\partial^{2}_{yy}V(t,y)+\sum_{i=1}^{n}\lambda^{i}\mathbb{E}\left[V(t,y+y\,\pi^{i}(e^{Z^{i}_{1}}-1))-V(t,y)\right]+\frac{(cy)^{\delta}}{\delta}\bigg\}\,$,
@@
with $V(T,y)=y^{\delta}/\delta$. This stochastic control problem does not admit an analytical solution but we can instead characterize its solution in terms of a PIDE, so as to assess the accuracy of the proposed Algorithms 1 and 2. Following Øksendal and Sulem (2007), we assume the value function is of the form $V(t,y)=A(t)y^{\delta}/\delta$. Then optimizing the HJB equation (31) leads to first order conditions showing that the optimal consumption rate $c^{*}$ is independent from the wealth $y$ and given by
$c^{*}(t,y)=A(t)^{1/(\delta-1)},$ (32)
and the optimal investment strategy is given by a vector of constants $\pi^{*}(t,y)=\pi^{*}:=(\pi^{*,1},\ldots,\pi^{*,n})^{\top}$ satisfying
$(\mu-r)^{\top}+\sigma\Sigma\sigma^{\top}\pi^{*}(\delta-1)+\sum_{i=1}^{n}\lambda^{i}\,\mathbb{E}\left[\left(1+\pi^{*,i}(e^{Z^{i}_{1}}-1)\right)^{\delta-1}\left(e^{Z^{i}_{1}}-1\right)\right]=0.$
@@
Plugging these optimal controls back into the HJB equation (31), we find that the function $A(t)$ satisfies the ODE
$\frac{\partial_{t}A(t)}{A(t)}=\rho\hskip 1.42262pt$- $\delta\left(r+(\mu-r)^{\top}\pi^{*}-A(t)^{1/(\delta-1)}\right)-\frac{1}{2}\delta(\delta-1)\pi^{*\top}\sigma\Sigma\sigma^{\top}\pi^{*}$ (33)
$\hskip 59.75078pt-\sum_{i=1}^{n}\lambda^{i}\mathbb{E}\left[\left(1+\pi^{*,i}(e^{Z^{i}_{1}}-1)\right)^{\delta}\hskip-1.42262pt-1\right]-A(t)^{1/(\delta-1)},$
@@
with $A(T)=1$, which can be solved numerically using again the Runge–Kutta method of order 5(4). For both Algorithms 1 and 2, we sample the time and space points independently from the uniform distributions $\mathcal{U}_{[0,T]}$ and $\mathcal{U}_{[0,y_{b}]}$, respectively. The parameters of the optimal investment problem below are as follows: $T=1,y_{b}=150,r=0.02,\rho=0.045,\delta=0.7,\lambda=0.45\cdot\mathbf{1}_{n},\mu_{Z}=0.25\cdot\mathbf{1}_{n},\sigma_{Z}=0.2\cdot\mathbf{1}_{n},\mu=0.032\cdot\mathbf{1}_{n},$ $\sigma=I_{n}$, $\Sigma=0.2\cdot\mathbf{1}_{n\times n}$ with $\text{diag}(\Sigma)=\mathbf{1}_{n}$. Note that, compared to the LQR problem in Section 5.1, the proportionality factor $\xi$ is now set to $10$. Moreover, since $A=\mathbb{R}_{+}^{n+1}$, $\sigma_{\alpha_{out}}$ is chosen to be the softplus activation function. Instead of the DGM architecture, we here train a classical feedforward neural network with 4 hidden layers, each of 128 neurons. Finally, the ${\rm MAE}_{V}$ and ${\rm MAE}_{\alpha}$ metrics are again defined as in Section 5.1 on a test set of size $M$, uniformly sampled from $0,1]\times[0,y_{b}]$ with $V$ and $\alpha^{*}$ obtained from the Runge–Kutta method described above.
@@
In Figure 11 we again compare ${\rm MAE}_{V}$ and runtime of GPI-PINN and GPI-CBU as functions of the number of epochs $k$ for $n=10$ stocks in the consumption-investment problem (30) with and without jumps. We again see that the residual-based GPI-PINN Algorithm 1 tends to be more stable and accurate for larger epochs, despite having a higher runtime. When accounting for jumps in the dynamics (29), GPI-PINN becomes numerically very time-consuming, even for a 10-dimensional control problem. This issue is even amplified compared with the LQR problem (see Figure 2(a)), as the jump size now depends on the control $\pi$ in the wealth dynamics (29). In contrast, GPI-CBU again handles these jumps far more efficiently.
@@
Figure 11: Comparison of ${\rm MAE}_{V}$ (blue) and runtime in seconds (green) of GPI-PINN (solid line) and GPI-CBU (dashed line) as functions of the number of epochs $k$ for the optimal consumption-investment problem without jumps (left) and with jumps (right) for $n=10$ stocks.